OpenStudy (anonymous):

If 14 out of 20 people like chocolate milk, what percent of the people do not like chocolate milk?

sammixboo (sammixboo):

You got choices?

OpenStudy (anonymous):

yes 5% 30% 40% 70%

sammixboo (sammixboo):

it's 30%, because 20 - 14 = 6, so: 20 ÷ 6 = .3 Then you take .3 and times it by 100%, and if you do that you'll get 30%

sammixboo (sammixboo):

I meant 6 ÷ 20 = .3